# SoK: Exploring the Potential of Large Language Models for Improving Digital Forensic Investigation Efficiency

### v1

@@ 3.3.1. Word to Vectors
Word embeddings, as introduced by Mikolov et al. (2013), entail precise and high-dimensional vector representations for words, particularly suited for extensive datasets comprising billions of text entries. The authors explored model architectures for word vectorisation, achieving substantial enhancements in accuracy while requiring lower computational resources and reduced training time (Mikolov et al., 2013). In the realm of LLMs, the primary objective is to generate new text based on the extensive dataset, on which it was trained. For this purpose, Vaswani et al. (2017) introduced the transformer model, aided by the word-to-vector model. This architecture incorporates a self-attention mechanism, as well as encoder and decoder processes, enabling the model to rapidly and simultaneously focus on pertinent information.
3.3.2. Transformer Models
@@
The transformer model initially aimed at machine translation, translating input words into another language, begins with word embedding, where inputs, termed as tokens, are vectorised. Recognising word order is achieved through positional encoding, with two main techniques: absolute and relative. Absolute positional encoding assigns unique vectors to each position, enhancing the model’s ability to recognise word placement and facilitate position-specific attention (Ke et al., 2021). Relative positional encoding, on the other hand, calculates the relative positions of words by introducing a bias term that quantifies distances between positions, improving the model’s capacity to understand word relationships within a sequence (Ke et al., 2021).
@@
Self-attention, a core mechanism within the transformer, calculates the relationship among words in a sentence, allowing the model to assess each word’s similarity to others and generate unique representations for each (Al-Rfou et al., 2019). The decoder mirrors the encoder’s steps but uses different weights, starting with positional encoding and computing self-attention values to identify the sentence’s initial translation word.
@@
This transformer process, leveraging stacked self-attention and unique positional encoding, has significantly advanced NLP tasks, including machine translation, text generation, and summarisation, by executing these processes in parallel and optimising weights for both encoder and decoder (Vaswani et al., 2017; Acheampong et al., 2021).
3.4. Specifically Trained LLMs
@@
The transformer model and the self-attention mechanism have paved the way for researchers to train language models on trillions of tokens with billions of parameters. Several LLMs have been trained and harnessed, each tailored with specific capabilities for diverse fields such as security, chemistry, engineering, medicine, business, tourism, and language-related applications. These models are employed in tasks ranging from detecting security threats, analysing data, and generating synthetic actions to teaching, code generation, structured query generation, planning, assisting in medical education, clinical decision-making, leveraging clinical settings, clinical validation, understanding general patterns and decision-making, bias detection, addressing ethical issues, language translations, question answering, information extraction, and business process automation, among others (Chang et al., 2024; Xuan-Quy et al., 2023; Tsai et al., 2023; Bran and Schwaller, 2023; Thirunavukarasu et al., 2023; Karabacak and Margetis, 2023; Eggmann et al., 2023; Vidgof et al., 2023; Carvalho and Ivanov, 2023; Hou et al., 2023; White et al., 2023; Odena et al., 2021; Glukhov et al., 2023; Moore et al., 2023; Caines et al., 2023; Bonner et al., 2023; Zhu et al., 2023; Rajkumar et al., 2022). The fine-tuning and retraining capabilities of LLMs empower them to be adapted to specific tasks or behaviours in a predefined manner. Fine-tuning involves taking an already trained language model and retraining its existing weights and bias values using a new dataset specific to a particular domain. This process allows the LLM to be customised and refined for tasks beyond its original training, enhancing its applicability in specific contexts (Ziegler et al., 2020). This process results in a new model that is more tailored and focused on the specified domain. In existing literature, it is frequently observed that LLMs are fine-tuned with a particular emphasis on engineering and research-related fields. This targeted fine-tuning ensures that the model is adept at handling tasks and generating content specifically relevant to the intricacies of these domains (Chang et al., 2024).
4. Capabilities of Large Language Models
@@ 4.1. Programming/Coding
The ability to generate source code within a specific context is a crucial proficiency inherent to a language model (Alon et al., 2020). Xu et al. (2022) conducted a systematic evaluation of six LLMs for code generation across twelve diverse programming languages. The benchmarking process involved the use of the HumanEval benchmark and an evaluation dataset designed to assess the functional correctness of programs generated by an LLM (Chen et al., 2021). Initially, the dataset was employed for benchmarking Code, a GPT language model fine-tuned on publicly available code sourced from GitHub (Chen et al., 2021). The Mostly Basic Programming Problems (MBPP) stands as another benchmark, comprising 974 programming tasks. It serves as a frequently employed evaluation dataset for LLMs specialising in code-related tasks (Wei et al., 2022). Several LLMs explicitly trained for code generation include Code LLaMA, CodeGen, StarCoder, PanGu-Coder, PanGu-Coder2, WizardCoder, InCoder 6B, CodeGen-Mono 16B, Code-Davinci-001, Code-Davinci-002, PaLM-Coder-540B, CodeT5+, InstructCodeT5+, GPT-4 with Reflexion, CodeGeeX, AlphaCode, Santa-Coder (Rozière et al., 2024; Nijkamp et al., 2022; Li et al., 2023a; Shen et al., 2023b; Luo et al., 2024; Zhou et al., 2022; Chowdhery et al., 2023; Wang et al., 2023c; Shinn et al., 2023; Zheng et al., 2023; Li et al., 2022; Allal et al., [n.d.]; Christopoulou et al., 2022; Chen et al., 2023c; Ye et al., 2023; Brown et al., 2020; Fried et al., 2023). A higher value for both HumanEval and MBPP indicates greater accuracy in code generation for a given task. For detailed information, refer to Table 1, which presents the counts for HumanEval and MBPP, along with the trained parameter size for each LLM. Even with the benchmarking rates from HumanEval, Liu et al. (2023d) raised questions about the accuracy of the functionality of the generated code. To address this concern, they proposed EvalPlus, an automated test generation engine, to verify the functional correctness of LLM-generated code.
@@
Table 1. Trained parameter count, Humaneval and MBPP count for specifically trained LLMs for code generation.
Model Parameters HumanEval MBPP
Code LLaMA $34B$
62.2
61.2
CodeGen $16.1B$
29.28
35.28
CodeGen-MONO $16.1B$
29.7
42.4
StarCoder $15.5B$
40.8
49.5
PanGu-Coder $2.6B$
27.78
23
PanGu-Coder2 $15B$
61.64
WizardCoder $15B$
57.3
51.8
Code-Davinci-001 (GPT3) $175B$
39.0
51.8
Code-Davinci-002 (GPT3.5) $175B$
47.0
58.1
PaLM-Coder $540B$
36.0
47.0
CodeT5+ $16B$
30.9
InstructCodeT5+ $16B$
35.0
GPT-4 with Reflexion $*1.76T*$
91.0
77.1
Sanata-Coder $1.1B$
18
35
AlphaCode $1.1B$
17.1
Codex-12B $12B$
28.81
code-cushman-001 $12B$
33.5
45.9
InCoder 6B $6.7B$
16.4
21.3
* Estimated parameter count as value is not officially released (Rizzo et al., 2024).
4.2. Vision Assistance
@@
Traditional vision assistant systems face limitations in image processing or recognition, as they are typically trained on fixed types of datasets. However, with the emergence of LLMs, this paradigm has shifted towards utilising raw text as the source of supervision (Radford et al., 2021a; Tewel et al., 2022). Research on visual recognition language models is experiencing exponential growth, with the number of models surpassing 1,500 in 2023 (Zhang et al., 2023b). Radford et al. (2021a) introduced a novel method called Contrastive Language-Image Pre-training (CLIP). This method is efficient and capable of performing a wide range of tasks during pre-training. It enables the training of a model to learn a shared representation space for both images and text, facilitating a deeper understanding of the relationships between the two modalities. Ramesh et al. (2021) proposes a model for text-to-image generation, capable of generating images as combinations derived from textual input or sentences. Moreover, with the model named Generating Images with Large Language Models (GILL), it becomes feasible to generate text, retrieve images, generate novel images, and interleave the results into coherent multimodal dialogues (Koh et al., 2023). VisionLLM is a framework leveraging LLMs for diverse vision tasks with unified language instruction, demonstrating generality and flexibility (Wang et al., 2023a). It incorporates a language-guided image tokeniser and an LLM-based task decoder, capable of handling open-ended tasks based on provided language instructions (Wang et al., 2023a).
@@
Visual instruction tuning leverages language-only models, such as GPT-4, to generate multimodal language-image instruction following data. This data is then utilised to instruction-tune large multimodal models, such as Large Language and Vision Assistant (LLaVA) (Liu et al., 2023c, b; Achiam et al., 2023). The open-source LLaVA project introduces an end-to-end trained model, integrating a vision encoder with an LLM. Notably, LLaVA showcases multimodal chat capabilities. LLaVa exhibits the capability to engage with images, providing detailed descriptions and responding to queries with a reported accuracy of 92.53% (Liu et al., 2023c). This demonstrates its effectiveness in understanding and generating contextually relevant information about visual content (Liu et al., 2023c). MiniGPT-4 is an open-source, powerful visual instruction-tuned LLM, and it demonstrates versatility by generating stories and poems inspired by provided images and teaching users how to cook based on visual cues from food photos. This showcases its ability to understand and respond creatively to diverse visual stimuli (Zhu et al., 2024).
@@
Position-enhanced Visual Instruction Tuning (PVIT) represents an extended version of Multimodal Large Language Models (MLLMs). It facilitates region-level encoding in an image, enabling the model to discern and identify information within specific regions (Chen et al., 2023a). Like PVIT, GPT-4 Region of Interest (ROI) is an instruction-tuned LLM capable of extracting information such as colour, shape, material, and action from a specified region (Zhang et al., 2024). This model enables users to interact with both language and drawing bounding boxes to indicate the area of interest within an image (Zhang et al., 2024). Other MLLMs, such as Visual ChatGPT, InternGPT, Flamingo, BLIP-2, and Kosmos, are noted in the literature for their capacity to assist users in visual-related information (Wu et al., 2023b; Liu et al., 2023a; Alayrac et al., 2022; Li et al., 2023b; Radford et al., 2021b; Zhao et al., 2023b; Peng et al., 2023).
@@
Video information is gaining prominence in vision assistance, and Zhao et al. (2023a) has introduced a novel approach to automatically narrate lengthy videos using LLMs. UniVL is another language pre-trained model designed for both multimodal understanding and generation. It is capable of retrieving a video segment based on text descriptions, generating captions for given video clips, segmenting a video according to a provided text input, and conducting multimodal sentiment analysis of a video segment (Luo et al., 2020). VidIL and MOV are additional MLLMs with similar capabilities, demonstrating proficiency in video classification and video-language operations such as video captioning, video question answering, video caption retrieval, and video future event prediction (Wang et al., 2022; Qian et al., 2022).
@@
These MLLMs adhere to a shared task set, encompassing visual question answering, visual captioning, visual common-sense reasoning, visual generation, multimodal affective computing, visual retrieval, vision-language navigation, multimodal machine translation, visual question generation, and visual dialoguing, as summarised in Table 2 (Kiros et al., 2014; Uppal et al., 2022).
Table 2. Capabilities of MLLMs trained for vision assistance
@@ 4.3. Conversation
Specific LLMs are trained explicitly for meaningful and coherent dialogues with humans. An example is Dialogue Generative Pre-trained Transformer (DialoGPT), a fine-tuned model trained on 174 million conversations from Reddit (Zhang et al., 2020). DialoGPT exhibits the ability to provide human-like answers in tested conversations (Zhang et al., 2020). Dettmers et al. (2023) introduced a fine-tuning mechanism for LLMs named Quantized Pre-trained Language Model into Low-Rank Adapters (QLoRA). This allows the fine-tuning of large-parameter LLMs with low training costs. They introduced Guanaco, a fine-tuned LLM with 65 billion parameters, which achieved a performance level of 99.3%. Falcon-180B and Falcon-40B represent another set of open-source LLMs with 180 billion and 40 billion parameters. These models are trained to communicate in multiple languages, allowing users to engage in conversations in languages other than English (Penedo et al., 2023). To evaluate the accuracy of human-like dialogue systems, Ou et al. (2023) proposed a dialogue evaluation benchmark named DialogBench, comprising 12 dialogue tasks to assess the capabilities of LLMs. In their evaluation, they assessed 28 pre-trained and instruction-tuned LLMs, demonstrating that GPT-4, ChatGPT, and KwaiYii-13B-Chat emerged as the top three models for conversations in domains related to daily life and professional knowledge.
@@
In DF chat conversations, the significance lies in facilitating non-technical investigators to elucidate terminologies and areas lacking understanding. This serves a dual purpose, acting as an interactive teacher to enhance comprehension in discussions (Scanlon et al., 2023a).
4.4. Prompt Engineering
@@
Achieving quality outputs from LLMs often relies on providing well-crafted, meaningful, and precise input queries, known as input prompts. However, even human-defined natural language instructions may not consistently yield the best results. Prompt engineering is a methodology that involves carefully defining and instructing LLMs to generate more accurate and desirable outputs. Through thoughtful refinement of input prompts, prompt engineering aims to enhance the performance and effectiveness of LLMs in generating outputs that align more closely with user expectations and requirements (White et al., 2023; Zhou et al., 2023). This plays a crucial role in biasing LLMs towards specific domains or topics, enabling a more targeted and nuanced response. By carefully crafting prompts, users can guide LLMs to delve deeper into the nuances of their queries, leading to more accurate and relevant outputs. This approach enhances the model’s responsiveness to specific areas of interest, allowing users to fine-tune and tailor their interactions with the LLM for more precise and meaningful outcomes. Prompt engineering with LLMs is employed across various sectors, including but not limited to medical, engineering, construction, and healthcare (B, 2022; Polak and Morgan, 2023).
@@
ChainForge is an open-source Graphical User Interface (GUI) tool developed specifically for prompt engineering and hypothesis testing derived from LLMs that can be used in the above-mentioned fields to generate accurate and swift outputs (Arawjo et al., 2023).
4.5. Autonomous Agents

### v2

@@ 2.3.1 Word to Vectors
Word embeddings, as introduced by Mikolov et al. [41], entail precise and high-dimensional vector representations for words, particularly suited for extensive datasets comprising billions of text entries. The authors explored model architectures for word vectorisation, achieving substantial improvements in accuracy while requiring fewer computational resources and reduced training time [41]. In the realm of LLMs, the primary objective is to generate new text based on the extensive dataset, on which it was trained. For this purpose, Vaswani et al. [42] introduced the transformer model, assisted by the word-to-vector model. This architecture incorporates a self-attention mechanism, as well as encoder and decoder processes, enabling the model to rapidly and simultaneously focus on pertinent information.
2.3.2 Transformer Models
@@
The transformer model initially aimed at machine translation, translating input words into another language, begins with word embedding, where inputs, termed as tokens, are vectorised. Recognising word order is achieved through positional encoding, with two main techniques: absolute and relative. Absolute positional encoding assigns unique vectors to each position, enhancing the model’s ability to recognise word placement and facilitate position-specific attention [43]. Relative positional encoding, on the other hand, calculates the relative positions of words by introducing a bias term that quantifies distances between positions, improving the model’s capacity to understand word relationships within a sequence [43].
@@
Self-attention, a core mechanism within the transformer, calculates the relationship among words in a sentence, allowing the model to assess each word’s similarity to others and generate unique representations for each [44]. The decoder mirrors the encoder’s steps but uses different weights, starting with positional encoding and computing self-attention values to identify the sentence’s initial translation word.
@@
This transformer process, leveraging stacked self-attention and unique positional encoding, has significantly advanced NLP tasks, including machine translation, text generation, and summarisation, by executing these processes in parallel and optimising weights for both encoder and decoder [42, 45].
2.4 Specifically Trained LLMs
@@
The transformer model and the self-attention mechanism have paved the way for researchers to train language models on trillions of tokens with billions of parameters. Several LLMs have been trained and harnessed, each tailored with specific capabilities for diverse fields such as security, chemistry, engineering, medicine, business, tourism, and language-related applications. These models are employed in tasks ranging from detecting security threats, analysing data, and generating synthetic actions to teaching, code generation, structured query generation, planning, assisting in medical education, clinical decision-making, leveraging clinical settings, clinical validation, understanding general patterns and decision-making, bias detection, addressing ethical issues, language translations, question answering, information extraction, and business process automation, among others [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. The fine-tuning and retraining capabilities of LLMs empower them to be adapted to specific tasks or behaviours in a predefined manner. Fine-tuning involves taking an already trained language model and retraining its existing weights and bias values using a new dataset specific to a particular domain. This process allows the LLM to be customised and refined for tasks beyond its original training, enhancing its applicability in specific contexts [64]. This process results in a new model that is more tailored and focused on the specified domain. In existing literature, it is frequently observed that LLMs are fine-tuned with a particular emphasis on engineering and research-related fields. This targeted fine-tuning ensures that the model is adept at handling tasks and generating content specifically relevant to the intricacies of these domains [46].
2.5 Large Action Models
Despite the primarily text-based operations common in LLMs, challenges persist in their ability to manipulate tasks and operations effectively. To address this limitation, recent efforts have introduced innovative concepts such as the Large Action Model (LAM) by the Rabbit research team [^1] https://www.rabbit.tech/research. In this approach, the base model is trained to comprehend sequences of human-provided actions and commands, enabling it to execute these actions accordingly. Similarly, Microsoft introduced the concept of Visualisation-of-Thought (VoT) aimed at integrating human cognitive abilities, specifically the creation of mental images, into the model [65]. Through this approach, it has been demonstrated that MLLMs excel in visual tasks, thereby enabling the extension of action capabilities within a LAM to any LLM. These advancements signify promising directions toward enhancing the practical applicability and versatility of language models across various domains.
3 Capabilities of Large Language Models
@@ 3.1 Programming/Coding
The ability to generate source code within a specific context is a crucial proficiency inherent to a language model [66]. Xu et al. [67] conducted a systematic evaluation of six LLMs for code generation across twelve diverse programming languages. The benchmarking process involved the use of the HumanEval benchmark and an evaluation dataset designed to assess the functional correctness of programs generated by an LLM [68]. Initially, the dataset was used to benchmark Code, a GPT language model fine-tuned with publicly available code sourced from GitHub [68]. The Mostly Basic Programming Problems (MBPP) stands as another benchmark, comprising 974 programming tasks. It serves as a frequently used evaluation dataset for LLMs specialising in code-related tasks [69]. Several LLMs explicitly trained for code generation include Code LLaMA, CodeGen, StarCoder, PanGu-Coder, PanGu-Coder2, WizardCoder, InCoder 6B, CodeGen-Mono 16B, Code-Davinci-001, Code-Davinci-002, PaLM-Coder-540B, CodeT5+, InstructCodeT5+, GPT-4 with Reflexion, CodeGeeX, AlphaCode, Santa-Coder [70, 71, 72, 73, 74, 75, 33, 76, 77, 78, 79, 80, 81, 82, 83, 4, 84]. A higher value for both HumanEval and MBPP indicates greater precision in code generation for a given task. For detailed information, refer to Table 1, which presents the counts for HumanEval and MBPP, along with the trained parameter size for each LLM. Even with the HumanEval benchmarking rates, Liu et al. [85] raised questions about the accuracy of the functionality of the generated code. To address this concern, they proposed EvalPlus, an automated test generation engine, to verify the functional correctness of the LLM-generated code.* Estimated parameter count as value is not officially released [86].
@@
Table 1: Trained parameter count, Humaneval and MBPP count for specifically trained LLMs for code generation.
{tblr}
hlines = 0.05em, hline1 = -0.16em, hline2 = -0.16em, hline3,20 = -0.08em, Model & Parameters HumanEval MBPP
Code LLaMA 34B 62.2 61.2
CodeGen 16.1B 29.28 35.28
CodeGen-MONO 16.1B 29.7 42.4
StarCoder 15.5B 40.8 49.5
PanGu-Coder 2.6B 27.78 23
PanGu-Coder2 15B 61.64 -
WizardCoder 15B 57.3 51.8
Code-Davinci-001 (GPT3) 175B 39.0 51.8
Code-Davinci-002 (GPT3.5) 175B 47.0 58.1
PaLM-Coder 540B 36.0 47.0
CodeT5+ 16B 30.9 -
InstructCodeT5+ 16B 35.0 -
GPT-4 with Reflexion 1.76T* 91.0 77.1
Sanata-Coder 1.1B 18 35
AlphaCode 1.1B 17.1 -
Codex-12B 12B 28.81 -
code-cushman-001 12B 33.5 45.9
InCoder 6B 6.7B 16.4 21.3
3.2 Vision Assistance
@@
Traditional vision assistant systems face limitations in image processing or recognition, as they are typically trained on fixed types of datasets. However, with the emergence of LLMs, this paradigm has shifted towards utilising raw text as the source of supervision [87, 88]. Research on visual recognition language models is experiencing exponential growth, with the number of models surpassing 1,500 in 2023 [89]. Radford et al. [87] introduced a novel method called Contrastive Language-Image Pre-training (CLIP). This method is efficient and capable of performing a wide range of tasks during pre-training. It enables the training of a model to learn a shared representation space for both images and text, facilitating a deeper understanding of the relationships between the two modalities. Ramesh et al. [90] proposes a model for text-to-image generation, capable of generating images as combinations derived from textual input or sentences. Moreover, with the model named Generating Images with Large Language Models (GILL), it becomes feasible to generate text, retrieve images, generate novel images, and interleave the results into coherent multimodal dialogues [91]. VisionLLM is a framework leveraging LLMs for diverse vision tasks with unified language instruction, demonstrating generality and flexibility [92]. It incorporates a language-guided image tokeniser and an LLM-based task decoder, capable of handling open-ended tasks based on provided language instructions [92].
@@
Visual instruction tuning leverages language-only models, such as GPT-4, to generate multimodal language-image instruction following data. This data is then utilised to instruction-tune large multimodal models, such as Large Language and Vision Assistant (LLaVA) [93, 94, 7]. The open-source LLaVA project introduces an end-to-end trained model, integrating a vision encoder with an LLM. Notably, LLaVA showcases multimodal chat capabilities. LLaVa exhibits the capability to engage with images, providing detailed descriptions and responding to queries with a reported accuracy of 92.53% [93]. This demonstrates its effectiveness in understanding and generating contextually relevant information about visual content [93]. MiniGPT-4 is an open-source, powerful visual instruction-tuned LLM, and it demonstrates versatility by generating stories and poems inspired by provided images and teaching users how to cook based on visual cues from food photos. This showcases its ability to understand and respond creatively to diverse visual stimuli [95].
@@
Position-enhanced Visual Instruction Tuning (PVIT) represents an extended version of Multimodal Large Language Models (MLLMs). It facilitates region-level encoding in an image, enabling the model to discern and identify information within specific regions [96]. Like PVIT, GPT-4 Region of Interest (ROI) is an instruction-tuned LLM capable of extracting information such as colour, shape, material, and action from a specified region [97]. This model enables users to interact with both language and drawing bounding boxes to indicate the area of interest within an image [97]. Other MLLMs, such as Visual ChatGPT, InternGPT, Flamingo, BLIP-2, and Kosmos, are noted in the literature for their capacity to assist users in visual-related information [98, 99, 100, 101, 102, 103, 104].
@@
Video information is gaining prominence in vision assistance, and Zhao et al. [105] has introduced a novel approach to automatically narrate lengthy videos using LLMs. UniVL is another language pre-trained model designed for both multimodal understanding and generation. It is capable of retrieving a video segment based on text descriptions, generating captions for given video clips, segmenting a video according to a provided text input, and conducting multimodal sentiment analysis of a video segment [106]. VidIL and MOV are additional MLLMs with similar capabilities, demonstrating proficiency in video classification and video-language operations such as video captioning, video question answering, video caption retrieval, and video future event prediction [107, 108].
@@
These MLLMs adhere to a shared task set, encompassing visual question answering, visual captioning, visual common-sense reasoning, visual generation, multimodal affective computing, visual retrieval, vision-language navigation, multimodal machine translation, visual question generation, and visual dialoguing, as summarised in Table 2 [109, 110].
Table 2: Capabilities of MLLMs trained for vision assistance
@@ 3.3 Conversation
Specific LLMs are trained explicitly for meaningful and coherent dialogues with humans. An example is Dialogue Generative Pre-trained Transformer (DialoGPT), a fine-tuned model trained on 174 million conversations from Reddit [111]. DialoGPT exhibits the ability to provide human-like answers in tested conversations [111]. Dettmers et al. [112] introduced a fine-tuning mechanism for LLMs named Quantised Pre-trained Language Model into Low-Rank Adapters (QLoRA). This allows the fine-tuning of large-parameter LLMs with low training costs. They introduced Guanaco, a fine-tuned LLM with 65 billion parameters, which achieved a performance level of 99.3%. Falcon-180B and Falcon-40B represent another set of open-source LLMs with 180 billion and 40 billion parameters. These models are trained to communicate in multiple languages, allowing users to engage in conversations in languages other than English [113]. To evaluate the accuracy of human-like dialogue systems, Ou et al. [114] proposed a dialogue evaluation benchmark named DialogBench, comprising 12 dialogue tasks to assess the capabilities of LLMs. In their evaluation, they assessed 28 pre-trained and instruction-tuned LLMs, demonstrating that GPT-4, ChatGPT, and KwaiYii-13B-Chat emerged as the top three models for conversations in domains related to daily life and professional knowledge.
@@
In DF chat conversations, the significance lies in facilitating non-technical investigators to elucidate terminologies and areas lacking understanding. This serves a dual purpose, acting as an interactive teacher to enhance comprehension in discussions [13].
3.4 Prompt Engineering
@@
Achieving quality outputs from LLMs often relies on providing well-crafted, meaningful, and precise input queries, known as input prompts. However, even human-defined natural language instructions may not consistently yield the best results. Prompt engineering is a methodology that involves carefully defining and instructing LLMs to generate more accurate and desirable outputs. Through thoughtful refinement of input prompts, prompt engineering aims to enhance the performance and effectiveness of LLMs in generating outputs that align more closely with user expectations and requirements [56, 115]. This plays a crucial role in biasing LLMs towards specific domains or topics, enabling a more targeted and nuanced response. By carefully crafting prompts, users can guide LLMs to delve deeper into the nuances of their queries, leading to more accurate and relevant outputs. This approach enhances the model’s responsiveness to specific areas of interest, allowing users to fine-tune and tailor their interactions with the LLM for more precise and meaningful outcomes. Prompt engineering with LLMs is employed across various sectors, including but not limited to medical, engineering, construction, and healthcare [116, 117].
@@
ChainForge is an open-source Graphical User Interface (GUI) tool developed specifically for prompt engineering and hypothesis testing derived from LLMs that can be used in the above-mentioned fields to generate accurate and swift outputs [118].
3.5 Autonomous Agents